A tank of diesel fuel (SG=.790) exerts a pressure of 21.40 psig. How many feet of diesel are there in the tank?
Answer #1

p=21.40 psig=1.47 bar

s.g.=0.790

density of diesel (d)=0.79 X 1000=790 kg/m^3

now, p=d*g*h

h=p/g*d=(1.47 X 10^5)/(9.81X790)=18.967m=62.211 ft